# AgentCgroup: Understanding and Controlling OS Resources of AI Agents

### v1

@@ Abstract.
AI agents are increasingly deployed in multi-tenant cloud environments, where they execute diverse tool calls within sandboxed containers, each call with distinct resource demands and rapid fluctuations. We present a systematic characterization of OS-level resource dynamics in sandboxed AI coding agents, analyzing 144 software engineering tasks from the SWE-rebench benchmark across two LLM models. Our measurements reveal that (1) OS-level execution (tool calls and container/agent initialization) accounts for 56–74% of end-to-end task latency; (2) memory, not CPU, is the concurrency bottleneck; (3) memory spikes are tool-call-driven with a 15.4 peak-to-average ratio; and (4) resource demands are highly unpredictable across tasks, runs, and models. Comparing these characteristics against serverless, microservice, and batch workloads, we identify three mismatches in existing resource controls: a granularity mismatch (container-level policies vs. tool-call-level dynamics), a responsiveness mismatch (user-space reaction vs. sub-second unpredictable bursts), and an adaptability mismatch (history-based prediction vs. non-deterministic stateful execution). We propose AgentCgroup, an eBPF-based resource controller that addresses these mismatches through hierarchical cgroup structures aligned with tool-call boundaries, in-kernel enforcement via sched_ext and memcg_bpf_ops, and runtime-adaptive policies driven by in-kernel monitoring. Preliminary evaluation demonstrates improved multi-tenant isolation and reduced resource waste.
@@ 1. Introduction
We propose AgentCgroup, an eBPF-based resource controller that addresses these mismatches through hierarchical cgroup v2 structures aligned with tool-call boundaries, in-kernel enforcement via sched_ext (Heo et al., 2023) and memcg_bpf_ops (Zhu, 2026) for microsecond-level reaction, and runtime-adaptive policies driven by in-kernel monitoring, with graceful degradation.
@@ 4.1. Granularity Mismatch
The mismatch has two dimensions: temporally, container-level policies cannot track tool-call-level dynamics; across resource types, CPU and memory are managed jointly despite being decoupled in agent workloads. On the temporal axis, cgroup v2 hard limits (memory.max) force a binary choice: setting to peak (4060 MB in the worst case) wastes 93% of allocated memory (needed only 2% of the time), while setting to average triggers OOM kills during 1–2 second tool bursts, destroying all accumulated agent state. Retry loops (§3) compound this, as progressive memory accumulation means a limit adequate for early iterations may trigger OOM by the fifth. Soft limits (memory.high) fare no better: reclaim cannot distinguish the 185 MB framework baseline (Node.js heap, V8 JIT cache) from tool subprocess memory, causing GC pressure that degrades LLM response parsing; moreover, a single container-level threshold cannot differentiate git status (13.5 MB) from pytest (518 MB P95). Kubernetes QoS classes (Kubernetes Community, 2024b) face both limitations: Guaranteed incurs comparable waste, BestEffort risks stateful agent termination, and Burstable still cannot set per-tool-call quotas; they also tie CPU and memory into a single class, despite memory bursts concentrated in tool calls (98.5%) while CPU bursts are more dispersed (55.3%), and CPU–memory correlation varying from 0.84 to 0.50 across tasks (§3). The large image footprint further precludes spawning additional containers for finer granularity (initialization consuming 31–48% of task time, §3), reinforcing the need for sub-container resource domains.
@@ 4.2. Responsiveness Mismatch
Agent resource bursts last only 1–2 seconds with change rates reaching several GB/s, a burst-silence pattern distinct from all traditional workload categories (Table 1). Moreover, their timing is unpredictable due to non-deterministic tool-call sequences. This combination of fast and unpredictable bursts means controllers need to react in real time at kernel speed. PSI-driven solutions (systemd-oomd (systemd Project, 2024), Meta oomd (Meta/Facebook, 2024)) monitor cgroup Pressure Stall Information and take action (kill the cgroup or trigger swap) when pressure exceeds a threshold. Their design assumes (a) a sufficient time window for decision-making after the pressure signal appears, and (b) that kill or migration is an acceptable degradation path. Both assumptions fail for agents: the full cycle from PSI signal generation to user-space daemon reception, decision, and cgroup control file write takes tens of milliseconds, by which time the burst has already passed or triggered kernel intervention. PSI is also a container-level aggregate that cannot attribute pressure to specific tool calls; it reports “container has pressure” but not “pytest caused it.” Kubernetes VPA (Kubernetes Community, 2024c) adjusts resources at Pod restart level (stable) or minute-level in-place resize (Kubernetes Community, 2024a) (alpha feature), orders of magnitude slower than second-scale tool bursts, making within-execution adjustment impossible. When slow reaction does lead to OOM, recovery is itself slow: restarting a multi-GB agent container incurs initialization overhead consuming 31–48% of typical task time (§3), amplifying the cost of every missed burst.
@@ 4.3. Adaptability Mismatch
This unpredictability is compounded by the high cost of the traditional fallback: kill-and-restart imposes a triple penalty on agent workloads. First, slow recovery: multi-GB agent container images make cold-start recovery (pulling layers, re-initializing the container and agent framework) consume 31–48% of total task time (§3), orders of magnitude longer than serverless cold starts. Second, lost state: an OOM kill destroys minutes of accumulated in-process LLM context, and unlike microservices with external state stores, this context cannot be checkpointed or migrated. Third, non-deterministic re-execution: re-running the same task follows an entirely different solution path (§3), so restart does not even guarantee convergence to the original solution. Any strategy that relies on termination as fallback, including Kubernetes eviction, systemd-oomd (systemd Project, 2024), and static OOM kill, incurs all three penalties simultaneously. Effective resource management should therefore adapt at runtime based on real-time observation, with graceful degradation (throttling, freezing) rather than termination.
@@ 5. AgentCgroupDesign and Implementation
AgentCgroupaddresses the three mismatches identified in §4: fine-grained resource domains match tool-call-level dynamics; in-kernel eBPF enforcement reacts at microsecond timescales; and runtime-adaptive policies replace historical prediction with in-kernel monitoring and graceful degradation.
Fine-grained resource domains. The granularity mismatch (§4) arises because existing controls set a single policy per container, while agent demands vary per tool call. AgentCgroup organizes resources using a hierarchical cgroup v2 structure where each agent workload maps to a cgroup node with tool calls as child nodes, enabling per-tool-call resource constraints while maintaining overall workload budgets. For recovery, AgentCgroup uses cgroup v2 lifecycle primitives: freezing subtrees when tool calls exceed soft limits, and atomically killing subtrees when termination is necessary.
@@
In-kernel enforcement. The responsiveness mismatch (§4) arises because user-space controllers react at millisecond-to-minute timescales, while agent bursts last 1–2 seconds with unpredictable timing. AgentCgroup executes control logic directly at kernel cgroup enforcement points via eBPF at the per-tool-call child cgroups described above, enabling microsecond-level reaction without user-kernel round trips. On CPU, AgentCgroup uses sched_ext (Heo et al., 2023) to maintain per-workload and per-tool-call metadata in BPF maps, prioritizing latency-sensitive tool calls with automatic fail-safe reversion on errors. On memory, AgentCgroup uses memcg_bpf_ops hooks (Zhu, 2026) to implement custom throttling delays when a tool call cgroup breaches its soft limit or hard limit. When memory pressure rises, AgentCgroup eBPF program applies graduated responses based on priority (throttling via memory.high delays, freezing via cgroup.freeze) rather than termination, preserving agent state.
Runtime-adaptive policies with graceful degradation. The adaptability mismatch (§4) arises because history-based prediction fails for non-deterministic agent workloads, and kill-restart destroys accumulated LLM context. AgentCgroup uses eBPF to trace process creation and memory allocation changes in-kernel, detecting tool-call boundaries and resource dynamics in real time to determine per tool call control behavior without user-space polling. A lightweight user-space daemon handles cgroup lifecycle management and policy configuration via shared BPF maps. We implement a proof-of-concept prototype of AgentCgroup in C using libbpf and BPF CO-RE (Linux Kernel Community, 2024b). The prototype extends Agentsight(Zheng et al., 2025) monitor and implements in-kernel memory enforcement via memcg_bpf_ops (Zhu, 2026) and CPU scheduling via sched_ext (Heo et al., 2023), and runs on Linux 6.15 with memcg_bpf_ops patches (currently under upstream review).
@@ 7. Conclusion and Future Work
We characterized AI coding agent tasks across two models, finding OS execution accounts for 56–74% of latency with a memory peak-to-average ratio of 15.4. Based on these findings, we presented AgentCgroup, an eBPF-based resource controller addressing granularity, responsiveness, and adaptability mismatches through tool-call-aligned cgroup domains in memory and CPU, microsecond-level reaction, and runtime-adaptive policies. Our current evaluation is limited to trace replay with a proof-of-concept prototype; the characterization covers one agent framework and one benchmark. Future work will evaluate AgentCgroup with live agent workloads at production scale, extend characterization to diverse agent frameworks and domains beyond coding, and explore fine-grained resource control across diverse container runtimes. Moreover, the current prototype mainly focuses on cpu and memory resource control; challenges such as initialization overhead, large container images, and retry-induced accumulation remain open and require optimizations across the sandbox runtime and orchestration stack.

### v2

@@ Abstract.
AI agents are increasingly deployed in multi-tenant cloud environments, where they execute diverse tool calls within sandboxed containers, each call with distinct resource demands and rapid fluctuations. We present a systematic characterization of OS-level resource dynamics in sandboxed AI coding agents, analyzing 144 software engineering tasks from the SWE-rebench benchmark across two LLM models. Our measurements reveal that (1) OS-level execution (tool calls and container/agent initialization) accounts for 56–74% of end-to-end task latency; (2) memory, not CPU, is the concurrency bottleneck; (3) memory spikes are tool-call-driven with a 15.4 peak-to-average ratio; and (4) resource demands are highly unpredictable across tasks, runs, and models. Comparing these characteristics against serverless, microservice, and batch workloads, we identify three mismatches in existing resource controls: a granularity mismatch (container-level policies vs. tool-call-level dynamics), a responsiveness mismatch (user-space reaction vs. sub-second unpredictable bursts), and an adaptability mismatch (history-based prediction vs. non-deterministic stateful execution). We propose AgentCgroup, an intent-driven eBPF-based resource controller that exploits agents’ ability to declare resource needs and reconstruct execution strategies, using hierarchical cgroup structures aligned with tool-call boundaries, in-kernel enforcement via sched_ext and memcg_bpf_ops, and runtime-adaptive policies. Preliminary evaluation demonstrates improved multi-tenant isolation and reduced resource waste. AgentCgroup is open-source at https://github.com/eunomia-bpf/agentcgroup.
@@ 1. Introduction
To address these mismatches, we observe that agent workloads have a key property absent in traditional workloads: agents can understand and adapt their own resource behavior. We exploit this in AgentCgroup, an intent-driven eBPF-based resource controller that combines hierarchical cgroup v2 structures aligned with tool-call boundaries, in-kernel enforcement via sched_ext (Heo et al., 2023) and memcg_bpf_ops (Zhu, 2026), and runtime-adaptive policies with graceful degradation (§5).
@@ 4.1. Granularity Mismatch
The mismatch has two dimensions: temporally, container-level policies cannot track tool-call-level dynamics; across resource types, CPU and memory are managed jointly despite being decoupled in agent workloads. On the temporal axis, cgroup v2 hard limits (memory.max) force a binary choice: setting to peak wastes over 90% of allocated memory (peak demand occurs less than 2% of the time), while setting to average triggers OOM kills during tool bursts, destroying accumulated agent state. Retry loops compound this, as progressive memory accumulation means a limit adequate for early iterations may trigger OOM by the fifth (§3). Soft limits (memory.high) fare no better: reclaim cannot distinguish the stable framework baseline (Node.js heap, V8 JIT cache) from tool subprocess memory, causing GC pressure on the agent runtime; moreover, a single container-level threshold cannot differentiate git status (13.5 MB) from pytest (518 MB P95). Kubernetes QoS classes (Kubernetes Community, 2024b) face both limitations: Guaranteed incurs comparable waste, BestEffort risks stateful agent termination, and Burstable still cannot set per-tool-call quotas; they also tie CPU and memory into a single class despite the weak and variable CPU–memory correlation observed in agent workloads (§3). Large image sizes further preclude spawning additional containers for finer granularity (initialization consuming 31–48% of task time), reinforcing the need for sub-container resource domains.
@@ 4.2. Responsiveness Mismatch
Agent resource bursts last only 1–2 seconds with change rates reaching several GB/s, a burst-silence pattern distinct from all traditional workload categories (Table 1). Moreover, their timing is unpredictable due to non-deterministic tool-call sequences. This combination of fast and unpredictable bursts means controllers need to react in real time at kernel speed. PSI-driven solutions (systemd-oomd (systemd Project, 2024), Meta oomd (Meta/Facebook, 2024)) monitor cgroup Pressure Stall Information and take action (kill the cgroup or trigger swap) when pressure exceeds a threshold. Their design assumes (a) a sufficient time window for decision-making after the pressure signal appears, and (b) that kill or migration is an acceptable degradation path. Both assumptions fail for agents: the full cycle from PSI signal generation to user-space daemon reception, decision, and cgroup control file write takes tens of milliseconds, by which time the burst has already passed or triggered kernel intervention. PSI is also a container-level aggregate that cannot attribute pressure to specific tool calls; it reports aggregate container-level pressure without attribution to individual tool calls. Kubernetes VPA (Kubernetes Community, 2024c) adjusts resources at Pod restart level (stable) or minute-level in-place resize (Kubernetes Community, 2024a) (alpha feature), orders of magnitude slower than second-scale tool bursts, making within-execution adjustment impossible. When slow reaction does lead to OOM, the high cost of agent container restart (§4) amplifies the cost of every missed burst.
@@ 4.3. Adaptability Mismatch
This unpredictability is compounded by the high cost of the traditional fallback: kill-and-restart imposes a triple penalty on agent workloads. First, slow recovery: cold-start of multi-GB containers consumes 31–48% of total task time (§3), orders of magnitude longer than serverless cold starts. Second, lost state: an OOM kill destroys minutes of accumulated in-process LLM context, and unlike microservices with external state stores, this context cannot be checkpointed or migrated. Third, non-deterministic re-execution: re-running the same task follows an entirely different solution path (§3), so restart does not guarantee convergence to the original solution. Any strategy that relies on termination as fallback incurs all three penalties. Effective resource management should therefore use graceful degradation (throttling, freezing) rather than termination, and exploit agents’ ability to adapt their behavior in response to resource feedback.
@@ 5. AgentCgroupDesign and Implementation
AgentCgroupaddresses the three mismatches identified in §4 with three corresponding mechanisms: fine-grained resource domains match tool-call-level dynamics (granularity); in-kernel eBPF enforcement at microsecond timescales (responsiveness); and intent-driven resource coordination replaces historical prediction with bidirectional adaptation (adaptability).
Fine-grained resource domains. To address the granularity mismatch (§4), AgentCgroup organizes resources using a hierarchical cgroup v2 structure where each agent workload maps to a cgroup node with tool calls as child nodes, enabling per-tool-call resource constraints while maintaining overall workload budgets. Concretely, a transparent bash wrapper intercepts each bash -c invocation, a common tool-call execution interface in agent frameworks, and creates an ephemeral child cgroup (tool_<pid>_<ts>/) for the duration of the command. The wrapper moves the tool process into this child cgroup, collects per-tool-call resource metrics (memory.peak, duration), and removes the cgroup upon exit. This requires no modification to any agent framework and adds negligible overhead per invocation. memcg_bpf_ops attached to the parent cgroup automatically applies to all child cgroups via cgroup v2 hierarchy inheritance, requiring no BPF code changes.
@@
In-kernel eBPF enforcement. To address the responsiveness mismatch (§4), AgentCgroup executes control logic directly at kernel cgroup enforcement points via eBPF, enabling microsecond-level reaction without user-kernel round trips. On CPU, AgentCgroup uses sched_ext (Heo et al., 2023) to maintain per-workload and per-tool-call metadata in BPF maps, prioritizing latency-sensitive tool calls with automatic fail-safe reversion on errors. On memory, AgentCgroup uses memcg_bpf_ops hooks (Zhu, 2026) to implement custom throttling delays when a tool call cgroup breaches its soft or hard limit. When memory pressure rises, the eBPF program applies graduated enforcement based on priority: throttling via memory.high delays, then freezing via cgroup.freeze, rather than termination, preserving accumulated agent state.
Intent-driven resource coordination. To address the adaptability mismatch (§4), AgentCgroup establishes a bidirectional protocol that exploits agents’ ability to adapt their resource behavior. Upward (agent system): before each tool call, the agent may declare its expected resource need via an environment variable (e.g., AGENT_RESOURCE_HINT="memory:high" for a test suite, "memory:low" for a file read); the wrapper maps these to per-tool-call memory.high limits. Downward (system agent): when a tool call is OOM-killed or throttled beyond recovery, the wrapper injects natural-language feedback into stderr (e.g., peak memory usage and a suggestion to reduce scope), enabling the agent to retry with a less resource-intensive approach. Declarations are advisory; the feedback loop corrects underestimates. When in-kernel enforcement alone cannot prevent a violation, this intent-driven channel provides a final graceful degradation path before termination. A lightweight user-space daemon manages cgroup lifecycle and policy configuration via shared BPF maps. We implement a proof-of-concept prototype in C using libbpf and BPF CO-RE (Linux Kernel Community, 2024b), extending Agentsight (Zheng et al., 2025). The prototype runs on Linux 6.15 with memcg_bpf_ops patches (currently under upstream review).
@@ 7. Conclusion and Future Work
We characterized AI coding agent tasks across two models, finding OS execution accounts for 56–74% of latency with a memory peak-to-average ratio of 15.4. Based on these findings, we presented AgentCgroup, an intent-driven eBPF-based resource controller addressing granularity, responsiveness, and adaptability mismatches through tool-call-aligned cgroup domains, microsecond-level in-kernel reaction, and resource adaptation that exploits agents’ ability to reconstruct execution strategies. Our current evaluation is limited to trace replay with a proof-of-concept prototype; the characterization covers one agent framework and one benchmark. Future work will evaluate AgentCgroup with live agent workloads at production scale, validate intent-driven adaptation across diverse tasks and agent frameworks, and explore fine-grained resource control across diverse container runtimes. Moreover, the current prototype mainly focuses on CPU and memory resource control; challenges such as initialization overhead, large container images, and retry-induced accumulation remain open and require optimizations across the sandbox runtime and orchestration stack.